For Questions 12 – 13, find the FICA and Medicare tax for each of the following self-employed people. Assume that none of them has earned over $127,200 for the year. Use a FICA rate of 12.4% and a Medicare rate of 2.9%.

12. Lois Steem is a self-employed assertiveness training coach. Her adjusted income is $19,500 for the first quarter. (a) FICA tax ________________________ (b) Medicare tax ________________________

13. The employees of the Hadley-Newham Funeral Home paid a total of $678.15 in Social Security tax, $158.60 in Medicare, and $2187.58 in federal withholding. Find the total amount the company should send the IRS.

SHOW ALL WORK!!!
Answer #1

Ans to Q.12

Adjusted Income of Lios Steem $19,500

Self Employed Deduction = 7.65%

Net Income = $19,500 - ($19,500 *7.65%) = $18,008.25

Social Security = $18,008.25 * 12.4% = $2,233

Medicare = $18,008.25 * 2.9% = $5,22

Total FICA    = $2,755

Ans to Q.13

Total amount the company should send the IRS :-

Social Security tax $678.15 (Employer's Share)

Medicare $158.60  (Employer's Share)

Federal withholding $2187.58

Total Amount   $3,024.33